Hi
Some sweets were distributed to a group of children. There were as many boys as there were girls. Each girl was given 5 sweets and each boy was given 8 sweets. The boys received 183 more sweets than the girls. How many children were there.

Found 3 solutions by mccravyedwin, greenestamps, math_tutor2020:
Answer by mccravyedwin(408)

Hi
Some sweets were distributed to a group of children. 
There were as many boys as there were girls. 

There were n boys and n girls.

Each girl was given 5 sweets 

So the girls had 5n sweets

and each boy was given 8 sweets. 

So the boys had 8n sweets.

The boys received 183 more sweets than the girls. 

So 8n is 183 more than 5n.

8n = 5n + 183

How many children were there?

Subtract 5n from each side

3n = 183

divide both sides by 3

n = 61

So there were 61 boys and 61 girls, or 122 children.

Edwin


Answer by greenestamps(13200)



The numbers of boys and girls were the same; each boy received 3 more sweets than each girl; and all together the boys received 183 more sweets than the girls. That means the number of boys was 183/3 = 61.

So there were 61 boys an 61 girls, for a total of 122.

ANSWER: 122


Answer by math_tutor2020(3817)


Each boy gets 8-5 = 3 extra sweets compared to each girl.
b = number of boys
3b = number of extra sweets the boys get compared to the girls.
3b = 183
b = 183/3
b = 61
61 boys + 61 girls = 122 children total.